OpenStudy (anonymous):

Use DeMoivre's formula to write cos(6X) in terms of sin(x) and cos(x)

OpenStudy (anonymous):

\[\Large \cos\alpha+i\sin\alpha=e^{i\alpha} \] De Moivre states that: \[\Large \left(\cos\alpha+i\sin\alpha \right)^n=\cos(n\alpha)+i\sin(n\alpha) \] So you can plugin \(n=6\) and then multiply out the binomial progression, compare coefficients afterwards.

OpenStudy (anonymous):

To be a bit clearer, compare real and imaginary parts with one another.